OpenStudy (anonymous):

write the equation in slope-intercept form and identify the slope and y intercept 7x+6y=17 slope intercept form is y=? the slope is ? y-intercept is?

OpenStudy (jdoe0001):

if you solve that for "y", what would you get?

OpenStudy (anonymous):

I need help wit the steps.and this website is moving soslow

OpenStudy (jdoe0001):

yeap, sadly this site gets like so

OpenStudy (anonymous):

I hate it. its like I have to send a message because it wont let me reply here

OpenStudy (jdoe0001):

well recall to solve linear equations you isolate the variable on one side so to isolate "y" on the left-hand-side subtract "7x" to both sides then divide by "6" both sides what would you get on that?

OpenStudy (anonymous):

so would that be 7x+6y-7x=17-7?

OpenStudy (jdoe0001):

well 7x to both sides you're subtracting it from the left, but on the right side you're only subtracting 7

OpenStudy (anonymous):

so 17-7x?

OpenStudy (anonymous):

so would it be 6y=10x

OpenStudy (jdoe0001):

on the right, yes, thus \(\bf \cancel{ 7x }+6y\cancel{ -7x }=17-7x\) now if you divide 6 on both?

OpenStudy (anonymous):

will it be 4? cause its not dividing with 10 so I just added 17+7 and got 24

OpenStudy (jdoe0001):

\(\large { \cancel{ 7x }+6y\cancel{ -7x }=17-7x\implies 6y=17-7x \\ \quad \\ \cfrac{\cancel{ 6 }y}{\cancel{ 6 }}=\cfrac{17-7x}{6}\implies y=\cfrac{17-7x}{6} \\ \quad \\ slope-intercept\ form\to y=mx+b\qquad thus \\ \quad \\ y=\cfrac{17-7x}{6}\implies y=\cfrac{17}{6}-\cfrac{7x}{6}\implies y=-\cfrac{7x}{6}+\cfrac{17}{6} \\ \quad \\ \implies \begin{array}{llll} y=&{\color{purple}{ -\cfrac{7}{6}}}x&{\color{blue}{ +\cfrac{17}{6}}}\\ &\quad \uparrow &\quad \uparrow \\ &{\color{purple}{ slope}}&{\color{blue}{ y-intercept}} \end{array} }\)